40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1611     Entry time: Wed May 20 01:53:48 2009
Author: rob, pete 
Type: Update 
Category: Locking 
Subject: violin mode filters in drstep_bang 

Recently the watch script was having difficulty grabbing a lock for more than a few seconds.  Rob discovered that the violin notch filters which were activated in the script were causing the instability.  We're not sure why yet.  The script seems significantly more stable with that step commented out.

ELOG V3.1.3-